3.618X10^24 molecules of sodium hydroxide (NaOH) would also be _____ grams NaOH
1
Expert's answer
2020-03-26T03:02:17-0400
"n = N\/N_a"

"n = 3.618*10^{24}\/6.022*10^{23} = 6.008 mol"

"m(NaOH) = n*M"

"n(NaOH) = 6.008*39,997 = 240.302g"

Answer: 240.302g